OpenStudy (anonymous):

Estimate the instantaneous rate of change of the volume after 5 hours in V(t)=t3-6t2+35

OpenStudy (bahrom7893):

V = t^3 - 6t^2 + 35 V' = 3t^2 - 12t V'(5) = 3*25 - 60 = 15